

********************************************************************
UARS WEEKLY STATUS REPORT - 18 June 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
During this past week, SUSIM successfully gathered its daily mid
(1.1 nm) resolution, MgII, occultation, solar pointing offset, and
continuous monitoring scans for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2000079. MEPS data files between
2000080 and 2000309 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 June 2004 (Orbit 69676, GMT Day 157) through 11
June 2004 (Orbit 69780, GMT Day 163).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 161/1231 69744 FULLTIME OPS
SUSIM 161/1452 69745 POWER ON
HALOE 163/1446 69775 POWER ON
A Reverse-to-Forward Yaw Around maneuver (number 130) was successfully
executed in Orbit 69700 beginning 158/ 15:06:29Z and completing at
158/ 15:43:39Z on 06 June 2004. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 55.71 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.08 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.62 and 20.04 percent
respectively. The next (forward-to-reverse) YAW is scheduled for
17 July 2004 (DOY 199).
The beta angle ranged between the angles of 05.5 degrees to 17.9
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 27 June 2004 (DOY 180).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.91 to 1.92 degrees C. Additional
battery performance data for each battery on 04 June and 11 June 2004
are provided in Appendix D.
The clock error ranged from 11.1 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
SOLSTICE = 01
SUSIM = 01
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT attempted to recreate the yaw wheel anomaly on the simulator
this week. Unfortunately, there were some problems getting the UTTS
scenario to work properly, so this work will continue next week.
A few days before the previous yaw maneuver high differential values
were seen on Battery 2, but no indication of that has been seen prior
to the yaw maneuver. The FOT will keep a close eye on the
differential voltage after the yaw.
MEETINGS
The FOT attended the UARS monthly DRB meeting on 08 June 2004 to
discuss the new Eclipse software delivery scheduled for next week.
The FOT also attended the biweekly meeting with the HRDI team to
discuss the analysis of the pointing issue. The analysis seems to
lean towards a possible timing error. More investigation will be done
into the FSW to determine if this is indeed the case.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 June 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69685 157/1514 04-161 MPS VT L6 (YAW)
69700 158/1435 04-161 MACS YAW START
69701 158/1559 04-161 MACS YAW COMPLETE
69701 158/1641 04-162 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69723 160/0256 04-163 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
69731 160/1634 04-161 MPS VT L5 (YAW)
69732 160/1733 04-161 MPS BATSOC (YAW)
69744 161/1231 04-164 SOLSTICE FULLTIME OPS
69745 161/1452 04-165 SUSIM POWER ON PART 1
69746 161/1549 04-165 SUSIM POWER ON PART 1 (CONT.)
69759 162/1324 04-165 SUSIM POWER ON COMPLETE
69775 163/1446 04-166 HALOE POWER ON PART 1
69778 163/1942 04-168 HALOE POWER ON COMPLETE
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 June 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 June 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
69701 158/1651 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
69723 160/0256 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 June 2004
04 June 2004 - GMT Day 156 Beta = 08.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.36/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +22.4/-33.6 +4.51/+3.87 +27.2/-12.8 25.4 1.01 19.9
3 +22.4/-33.6 +3.55/+2.60 +26.4/-12.8 25.4 1.00 19.5
11 June 2004 - GMT Day 163 Beta = 17.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-089.6 +5.47/+3.55 +28.0/-14.8 24.5 1.03 19.5
3 +00.0/-123.2 +4.51/+2.60 +26.8/-14.4 24.5 1.02 18.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
69685 157/1514 LV5 -> LV6
69731 160/1634 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 June 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT 69775 163/1403 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-031 FOT 69662 156/0159 36027 TDRSS WEST OUTAGE
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT See List June 04 30832 Generic Late Acquisition